### v2

# Extracting Phonon Quasiparticles from Molecular Dynamics Simulations

Abstract
Phonon anharmonicity is ubiquitous in real materials and is crucial for understanding thermal properties and phase stability. In this work, we show that phonon quasiparticles are optimally described by modes with maximum lifetimes, and prove that information about these quasiparticles is contained in two small matrices $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from molecular dynamics simulations. Based on these knowledge, we proposed an optimization scheme, which allows us to efficiently determine temperature-dependent phonon modes along with their frequencies and lifetimes. We verified this method by applying it to silicon and cubic CaSiO3, where it successfully captured their temperature-dependent phonon behaviors and the well-known phonon softening in cubic CaSiO3. This theory provides a convenient tool for investigating phonon quasiparticles and can be extended to study other quasiparticles, such as electrons, holes, and magnons.
Phonons are quantized vibrational modes in crystalline solids, describing how atoms collectively oscillate around their equilibrium positions. These fundamental quasiparticles are essential in physics textbooks, playing a crucial role in our understanding of heat transfer, superconductivity, and thermoelectric effects. Modern theoretical studies determine phonon modes and frequencies through harmonic approximation, using density functional theory (DFT) with either perturbation theory (DFPT)[1, 2, 3, 4] or the finite displacement method[5, 6, 7, 8]. The resulting phonons all have infinite life time. However, in real materials, phonons scatter with each other and with other quasiparticles, resulting in finite lifetimes and modified material properties.
Over the past decades, researchers have developed various methods to study phonons beyond the harmonic approximation. A natural method is to include phonon-phonon collisions using the third-order or even fourth-order force constants [9, 10, 11, 12, 13, 14, 15, 16]. This approach has effectively determined phonon lifetimes across various materials, including semiconductors [17, 18, 19, 20], rocksalt materials[21], zincblende compounds[22, 21, 23, 24] and wurtzite compounds [24], and two-dimensional crystals [25, 26, 27, 28]. Although these perturbative approaches have proven successful, numerous non-perturbative methods have also emerged, such as the self-consistent harmonic approximation (SCHA)[29, 30, 31, 32], self-consistent ab initio lattice dynamics (SCAILD) [33], stochastic self-consistent harmonic approximation (SSCHA)[34], self-consistent phonon calculations (SCPH) [35], Wigner transport equation (WTE)[36, 37] and quantum self-consistent ab initio lattice dynamics (QSCHILD) [38]. Additionally, molecular dynamics (MD), especially the ab initio molecular dynamics (AIMD), offers another powerful non-perturbative approach for studying phonon anharmonicity, as atomic trajectories naturally incorporate anharmonic effects to all orders. The temperature-dependent effective potential (TDEP) method[39] extracts effective force constants from AIMD simulations, enabling the calculation of anharmonic phonon properties. Similarly, the phonon quasiparticle approach (PQA) projects AIMD trajectories onto normal modes directly obtain anharmonic phonon frequencies and lifetimes[40, 41, 42]. With all these methods yielding phonon quasiparticles via distinct approaches, a fundamental question arises: How accurate are these phonon quasiparticles in modeling the vibration of a real system?
The key is lifetime. In fact, phonon quasiparticles with inaccurate phonon modes typically exhibit shorter lifetimes. Therefore, modes with maximized lifetimes provide the most accurate representation of of phonon quasiparticles. This principle inspires a concise and efficient method, developed in this study, to extract phonon quasiparticles from MD simulations.
For a pure harmonic system with $n$ atoms in primitive cell, the $j$-th phonon mode at $\vec{k}$ point of Brillouin zone is $\varphi_{i\sigma}^{\vec{k},j}$ ($i=1,2,..,n;\,\sigma=x,y,z$). There are 3nn phonon modes in total, forming a complete orthonormal basis set. Then, the atom trajectories or velocities in a MD simulation can be expanded as $\mathcal{R}_{i\sigma}^{I}(t)=Re\left(\sum_{\vec{k},j}c^{\vec{k},j}\varphi_{i\sigma}^{\vec{k},j}e^{\mathbf{i}(\vec{k}\cdot\vec{R}_{I}-\omega_{\vec{k},j}t)}\right)$, where $\omega$ is eigen frequency, $t$ is time, and $\vec{R}_{I}$ is the displacement vector of the $I$-th cell. The coefficient $c_{\vec{k},j}$ included here represents the weight of mode $\varphi^{\vec{k},j}$. Since $\mathcal{R}_{i\sigma}^{I}(t)$ contains phonon modes at all $\vec{k}$ points, we can filter out the modes at specific $\vec{k}$ by real space Fourier transformation:
$f^{\vec{k}}_{i\sigma}(t)=\frac{1}{N}\sum_{I}\mathcal{R}_{i\sigma}^{I}(t)e^{-\mathbf{i}\vec{k}\cdot\vec{R}_{I}}$ (1)
$=\frac{1}{2}\sum_{j}c^{\vec{k},j}\varphi_{i\sigma}^{\vec{k},j}e^{-\mathbf{i}\omega_{\vec{k},j}t}+\frac{1}{2}\sum_{j}(c^{-\vec{k},j})^{*}(\varphi_{i\sigma}^{-\vec{k},j})^{*}e^{\mathbf{i}\omega_{-\vec{k},j}t}$
It turns out that the result contains both positive frequencies from $\vec{k}$ point and negative frequencies from $\vec{k}$ point. To isolate contributions solely from the $\vec{k}$ point, we perform time Fourier transformation, remove negative frequencies, and apply inverse Fourier transformation, yielding:
$g^{\vec{k}}(t)=\frac{1}{2}\sum_{j}c^{\vec{k},j}\varphi^{\vec{k},j}e^{-\mathbf{i}\omega_{\vec{k},j}t}$ (2)
We have omitted the subscripts $i$ and $\sigma$ for simplicity, which makes $\varphi^{\vec{k},j}$ and $g^{\vec{k}}(t)$ column vectors with 3nn elements.
Eq. 2 also applies to real systems. The key differences are that both modes and frequencies depend on MD temperature, and the coefficients $c^{\vec{k},j}$ become time-dependent due to mode scattering caused by anharmonicity. A non-zero $|\frac{dc^{\vec{k},j}(t)}{dt}|$ leads to a finite quasiparticle lifetime, which further depends on the choose of modes. If inaccurate modes and the associated frequencies are used to expand MD trajectories or velocities, the coefficients $c^{\vec{k},j}(t)$ will vary more rapidly over time. This suggests that optimal phonon quasiparticles can be defined as those modes that minimize $|\frac{dc^{\vec{k},j}(t)}{dt}|$, equivalently speaking maximize quasiparticle lifetimes.
Next, we describe the method for obtaining optimal phonon quasiparticles. After a MD simulation, the column vector $g^{\vec{k}}(t)$ is constructed directly from atom trajectories, velocities, or their mass-weighted forms. The time-dependent coefficient is then given by $c^{\vec{k},j}(t)=2(\varphi^{\vec{k},j})^{+}g^{\vec{k}}(t)e^{\mathbf{i}\omega_{\vec{k},j}t}$. For mathematical convenience, we minimize $|\frac{dc^{j}(t)}{dt}|^{2}$ rather than $|\frac{dc^{\vec{k},j}(t)}{dt}|$ to optimize phonon quasiparticles. Since the optimization was performed individually for each $\vec{k}$ point, for brevity, we omit the $\vec{k}$ superscript in the following discussions, unless needed for clarity.
Figure 1: Optimization process and ACF for silicon phonons. (a) $\Delta C_{v}$ obtained using mass-weighted velocities for different initial modes. (b) $\Delta C_{r}$ obtained using mass-weighted trajectories. In both panels, $\Delta C$ represents the difference between $C$ and its optimal value. (c) Normalized ACF of an optical phonon. (d) Normalized ACF of an acoustic phonon. All calculations performed at $\vec{k}=(1/5,1/5,1/5)$ at 500K.
Starting from an initial guess of the mode set $\varphi^{j}$ (jj=1,2,…,3nn), we compute the sum of $|\frac{dc^{j}(t)}{dt}|^{2}$ and then take its time average:
$\begin{aligned} C=\frac{1}{t_{f}-t_{i}}\sum_{j}\int_{t_{i}}^{t_{f}}|\frac{dc^{j}(t)}{dt}|^{2}dt=\sum_{j}(\varphi^{j})^{+}\mathcal{G}^{j}\varphi^{j}\end{aligned},$
where $t_{i}$ and $t_{f}$ are the initial and final simulation times, and $\mathcal{G}^{j}=\mathcal{G}_{0}+\omega_{j}\mathcal{P}+\omega_{j}^{2}\mathcal{Q}$. The matrices $\mathcal{G}_{0}$, $\mathcal{P}$, and $\mathcal{Q}$ here are independent of the initial guessed phonon modes. They are all Hermitian and can be constructed solely from the MD trajectory,
$\mathcal{G}_{0}$ $=\frac{4}{t_{f}-t_{i}}\int_{t_{i}}^{t_{f}}\frac{dg(t)}{dt}(\frac{dg(t)}{dt})^{+}dt\,$,
$\mathcal{Q}$ $=\frac{4}{t_{f}-t_{i}}\int_{t_{i}}^{t_{f}}g(t)(g(t))^{+}dt\,$,
$\mathcal{P}$ $=\mathbf{i}[(\mathcal{S})^{+}-\mathcal{S}]\,$,
$\mathcal{S}$ $=\frac{4}{t_{f}-t_{i}}\int_{t_{i}}^{t_{f}}\frac{dg(t)}{dt}(g(t))^{+}dt\,.$
Once these matrices are constructed, the minimization of $C$ is straightforward. The optimal frequencies satisfy the relation $\frac{\partial C}{\partial\omega_{j}}=0$. Then we get
$\omega_{j}=-\frac{P_{jj}}{2Q_{jj}}=-\frac{Im(S_{jj})}{Q_{jj}},$ (3)
where $Im(S_{jj})$ pickes the imaginary part of $S_{jj}$, the matrix elements $P_{ij}=(\varphi^{i})^{+}\mathcal{P}\varphi^{j}$, $Q_{ij}=(\varphi^{i})^{+}\mathcal{Q}\varphi^{j}$, and $S_{ij}=(\varphi^{i})^{+}\mathcal{S}\varphi^{j}$.
The optimization of phonon modes requires a more complex process involving a series of rotations in Hilbert space, $\left[\varphi^{1},\varphi^{2},...,\varphi^{3n}\right]\rightarrow\left[\varphi^{1},\varphi^{2},...,\varphi^{3n}\right]U$, where the matrix $U$ is unitary. This rotation results in a variation of $C$:
$\begin{aligned} \delta C=\delta Tr\left(PU\Omega U^{+}\right)+\delta Tr\left(QU\Omega^{2}U^{+}\right)\end{aligned}.$
The $Tr$ notation here represents the trace of a matrix. The matrix $\Omega$ is diagonal, with its diagonal elements corresponding to the mode frequencies. As we have shown, these frequencies depend on the phonon modes, and are therefore dependent on $U$. This variation equation leads us to:
$\begin{aligned} \frac{\delta C}{(\delta U)_{ij}}=(\Omega_{jj}-\Omega_{ii})P_{ji}+(\Omega_{jj}^{2}-\Omega_{ii}^{2})Q_{ji}\end{aligned}.$
The derivation of these two equations is somewhat tedious, we present them in the supplementary materials for completeness.
Let’s define a matrix $B$, whose elements are
$B_{ij}$ $=\left(\frac{\delta C}{\delta U_{ij}}\right)^{*}=(\Omega_{jj}-\Omega_{ii})P_{ij}+(\Omega_{jj}^{2}-\Omega_{ii}^{2})Q_{ij}$ (4)
$=\mathbf{i}(\Omega_{jj}-\Omega_{ii})(S_{ji}^{*}-S_{ij})+(\Omega_{jj}^{2}-\Omega_{ii}^{2})Q_{ij}.$
This represents the fastest direction for increasing $C$. The complex conjugate operation ensures that $\delta C$ remains real. To minimize $C$, we define the unitary transformation as $U=e^{-\gamma B}$, where $\gamma$ is the step length of rotation in Hilbert space. This value should be small enough to maintain stability in the minimization process. In practical applications, we can use the approximation $U=1-\gamma B$. By repeatedly performing these calculations until $B$ is zero, we ultimately obtain the optimal phonon modes and their frequencies.
Having established this theoretical framework, we apply the optimization method to a real material, silicon. To enhance computational efficiency, we performed a $\times5$$\times$5 supercell MD simulation at 500 K and 1000 K over 50 ps using the first-principles machine learning force field in VASP code[43]. The simulation begins with a 4 ps canonical (NVT) ensemble MD to achieve thermal equilibrium, followed by a transition to the microcanonical (NVE) ensemble MD. The NVE ensemble is suitable here because it eliminates artificial interactions between phonon modes and the heat bath. Subsequently, function $g(t)$ was constructed based on the mass-weighted trajectories and velocities. Six phonon modes were randomly generated and orthogonalized using the Gram-Schmidt procedure. As shown in Fig. 1(a), the initial value of $\Delta C$ for $\vec{k}=(1/5,1/5,1/5)$ at 500 K based on mass-weighted trajectories is substantially large but decreased rapidly during optimization. The value of $\Delta C$ based on mass-weighted velocities has the same trend. These results demonstrate the high efficiency of the optimization method.
The efficiency of the optimization process can be further enhanced by selecting better initial modes. The harmonic phonon modes are a viable option. Additionally, the peak values from the Fourier transformation of Eq. 1 serve as effective initial modes. In fact, the eigenvectors of $\mathcal{G}_{0}$, $\mathcal{P}$, and $\mathcal{Q}$ provide even more convenient alternatives. Take $\mathcal{Q}$ as an example, combining its definition and the expression of $g(t)$ yields,
$\mathcal{Q}$ $=\frac{1}{t_{f}-t_{i}}\sum_{i,j}\varphi^{i}(\varphi^{j})^{+}\int_{t_{i}}^{t_{f}}c^{i}(t)(c^{j}(t))^{*}e^{\mathbf{i}(\omega_{j}-\omega_{i})t}dt$
$\approx\frac{1}{t_{f}-t_{i}}\sum_{i,j}\varphi^{i}(\varphi^{j})^{+}\delta_{\omega_{i},\omega_{j}}\int_{t_{i}}^{t_{f}}c^{i}(t)(c^{j}(t))^{*}dt.$
The integration in the first line consists of a slowly varying function $c^{i}(t)(c^{j}(t))^{*}$ and an exponential function $e^{\mathbf{i}(\omega_{j}-\omega_{i})t}$. The latter oscillates quickly if $\omega_{j}\neq\omega_{i}$, making the integration negligible and producing the delta function in the last line. This form indicates that the eigenvectors of $\mathcal{Q}$ are isolated anharmonic phonon modes and the linear combinations of degenerated anharmonic phonon modes, while the eigenvalues are their weights. The same argument applies to $\mathcal{G}_{0}$ and $\mathcal{P}$ as well, though with different eigenvalues. We should caution here that due to the finite MD simulation time and the phonon anharmonicity, the last line is only an approximation. As a result, the eigenvectors do not precisely represent the anharmonic phonon modes, which requires optimization. The optimization results using the eigenvectors of $\mathcal{G}_{0}$, $\mathcal{P}$, and $\mathcal{Q}$ as initial modes are displayed as red lines in Fig. 1(a,b). We can see that the initial values of $\Delta C$ are three orders of magnitude smaller than those obtained from random initial modes. Therefore, these initial modes are very effective, allowing us to get more accurate anharmonic phonons within the same number of optimization steps.
Figure 2: Phonon dispersions for silicon at (a) 500 K and (b) 1000 K. The linewidths are represented by the thickness of the curves. The corresponding phonon spectral functions are presented in panels (c) and (d).
We now proceed to calculate the phonon linewidth, which is inversely proportional to the phonon lifetime. Previous studies have shown that phonon linewidth can be obtained from the mode-projected velocity autocorrelation function (ACF)[41, 42]. Similarly, in this work, phonon linewidth can also be derived from the ACF of mode-projected $g(t)$ function, $g_{j}(t)=(\varphi^{j})^{+}g(t)$:
$<g_{j}(0)\cdot g_{j}(t)>$ $=\frac{1}{t_{f}-t_{i}}\int_{t_{i}}^{t_{f}}g_{j}^{*}(\tau)g_{j}(t+\tau)d\tau$ (5)
$=Q_{jj}e^{-i\omega_{j}t}e^{-\Gamma_{j}t}.$
The frequency and lifetime can be obtained from the time derivative of the ACF:
$-i\omega_{j}-\Gamma_{j}=\frac{<g_{j}(0)\cdot\dot{g}_{j}(0^{+})>}{Q_{jj}}=\frac{S_{jj}}{Q_{jj}}.$ (6)
Eq. 3 represents the imaginary part of this equation. For the real part, the value of $S_{jj}$ must be constructed using $\dot{g}_{j}$ at a time slightly later than $g_{j}$. This is because the exponential function in Eq. 5 is only valid for $t>0$. This requirement increases the calculation complexity, making it simpler to obtain phonon lifetime by fitting the ACF using Eq. 5.
The central equations of this method, Eqs. 3 and 4, rely on the matrices $\mathcal{S}$ and $\mathcal{Q}$, which are constructed directly from molecular MD trajectories or velocities. The procedure begins by diagonalizing $\mathcal{Q}$ to obtain its eigenvectors, which serve as the initial phonon modes. These modes are then optimized by constructing the matrices $S$ and $Q$ and applying Eqs. 3 and 4, yielding the optimal phonon modes and frequencies. Finally, the linewidths are calculated from ACF.
The autocorrelation function for silicon are shown in Fig. 1(c,d). Panel (c) displays an optical phonon with a frequency of 13.9 THz and lifetime of 0.8 ps. Panel (d) shows an acoustic phonon with a frequency of 2.3 THz and a significantly longer lifetime of 36.1 ps.
For each $\vec{k}$ point, these equations enable us to determine the optimal phonon modes, frequencies, and their linewidths. Let’s proceed to calculate the phonon dispersions. The dynamical matrix can be constructed as $D_{\vec{k}}=\varPhi_{\vec{k}}\widetilde{\Omega}_{\vec{k}}^{2}\Phi_{\vec{k}}^{+}$, where $\varPhi_{\vec{k}}=[\varphi^{\vec{k},1},\varphi^{\vec{k},2},...,\varphi^{\vec{k},3n}]$, and $\widetilde{\Omega}_{\vec{k}}$ is a complex matrix with elements $\widetilde{\omega}^{j}_{\vec{k}}=\omega^{j}_{\vec{k}}-\mathbf{i}\Gamma^{j}_{\vec{k}}$. The force constant matrix is given by $F_{\vec{R}}=\frac{1}{N}\sum_{\vec{k}}D_{\vec{k}}e^{\mathbb{i}\vec{k}\cdot\vec{R}}$, with the sum performed over all $\vec{k}$ points in the first Brillouin zone compatible with the supercell. Then the dynamical matrix at $\vec{k}^{\prime}$ points along the high-symmetry line of the first-Brillouin zone is $D_{\vec{k}^{\prime}}=\sum_{\vec{R}}F_{\vec{R}}e^{-\mathbb{i}\vec{k}^{\prime}\cdot\vec{R}}$. Finally, diagonalizing these matrices produces the phonon dispersions, linewidths, and modes.
The calculated phonon dispersions for silicon are shown in Fig. 2(a-b). MD simulations were performed at both 500 K and 1000 K, with the resulting phonon dispersions displayed in Fig. 2(a) and Fig. 2(b), respectively. These dispersions closely match the harmonic phonon dispersions, indicating that anharmonic effects are weak in this system. The three optical phonons display larger linewidths than the acoustic phonons. Due to larger vibration amplitudes causing stronger anharmonic effects, the phonon linewidths at 1000 K are significantly larger than those at 500 K. Furthermore, compared to the phonon dispersions at 500 K, all phonon frequencies at 1000 K are reduced, agree with previous theoretical[44] and experimental[45] studies. For instance, the highest frequency at the $\Gamma$ point is 14.5 THz at 500 K, which decreases to about 14.2 THz at 1000 K.
With the phonon frequencies and linewidths calculated along the high-symmetry lines, we can now plot the phonon spectral functions, which can be directly compared with experimental results. Based on the Bose-Einstein distribution and the standard Lorentzian function of spectral lines, the phonon spectral function at $\vec{k}$ point can be written as
$W(\omega)=\frac{1}{\pi}\sum_{j}\frac{1}{e^{\hbar\omega_{j}/kT}-1}\frac{\Gamma_{j}}{\hbar(\omega-\omega_{j})^{2}+\Gamma^{2}},$
where $\hbar$ is the reduced Planck constant, $k$ is the Boltzmann constant, and $T$ is temperature. The calculated results are shown in Fig. 2(c,d). At 500 K, low-frequency phonons were more excited, and the spectral functions of optical phonons are weak. In contrast, at 1000 K, high-frequency phonons were also excited, which gives the optical phonons strong intensity. The calculated spectrum agrees with the experimental results[46], except for the intensity of the highest optical phonon peak, which is significantly stronger in the experiment. This can be attributed to the optical phonon excitation triggered by neutral scattering.
Figure 3: Phonon dispersions for CaSiO3 at (a)400 K and (b) 1000 K. The linewidths are represented by the thickness of the curves. Panels (c) and (d) display the corresponding phonon spectral functions.
Let’s continue to study a more complex material, the perovskite CaSiO3 at high pressure. It has strong anharmonic effect of phonons[42, 47, 48], and is a main component of pyrolitic lower mantle[49, 50]. The harmonic phonon dispersions of cubic CaSiO3 show imaginary frequencies at all pressures, therefore, it is unstable and can change to more stable phases such as tetragonal or orthorhombic structures. Experiments have shown that the cubic CaSiO3 can be stabilized by high teperature and pressure[51, 52, 53]. Many researchers have studied its anharmonic phonons, and the study of this material is still active[54, 55, 56] in geoscience.
Unlike silicon, cubic CaSiO3 is a typical polar material. In such systems, the transverse optical (TO) phonon modes and longitudinal optical (LO) phonon modes split due to long-range Coulomb interactions. Therefore, after obtaining the effective force constants, we need to apply corrections[57, 58] based on Born effective charges and the dielectric tensor. The lattice parameter of the unit cell was set to 3.45 Å, corresponding to the pressure at the top of the lower mantle. The calculated phonon dispersions of cubic CaSiO3 at 400 K and 1000 K are shown in Fig. 3(a,b). The LO and TO phonons split at the $\Gamma$ point. The lowest three branches of phonons around the $R$ point have zero frequencies, corresponding to the imaginary frequencies in previous studies—indicating that cubic CaSiO3 is unstable at 400 K. However, these phonons have positive frequencies at 1000 K, showing that the material becomes stable at higher temperatures. The top three phonon branches throughout the entire Brillouin zone and the lowest three phonon branches at the $R$ point exhibit relatively large linewidths. The corresponding phonon spectral functions are shown in Fig. 3(c,d).
The effectiveness of this method stems from the fact that phonon quasiparticles govern atomic motion in crystals, with key information stored in compact matrices. This paradigm is broadly applicable to other systems governed by their own quasiparticles. For example, in the case of magnons, compact matrices can be derived from spin trajectories calculated using the Landau-Lifshitz-Gilbert equation. Similarly, for electrons and holes, compact matrices or kernel functions can be extracted from time-dependent DFT calculation results. The corresponding quasiparticles can then be systematically determined from these matrices or kernel functions.
In summary, this work defines optimal phonon quasiparticles as those that maximize their lifetimes. We demonstrate that all information about these quasiparticles is stored in two simple matrices, $\mathcal{S}$ and $\mathcal{Q}$, which can be constructed directly from MD trajectories or velocities. The eigen vectors of $\mathcal{Q}$ can be used as the initial phonon modes. Then the optimal phonons and their corresponding frequencies and lifetimes can be obtained from an optimization scheme. Tests on silicon show weak anharmonicity with temperature-dependent shifts, while cubic CaSiO3 exhibits strong anharmonic effects, being unstable at 400 K but stabilized at 1000 K. These tests indicate the reliability of this method. This theory and optimization method provide a clear picture of phonon quasiparticles in real materials, offering a convenient tool for their investigation. Its core principles are readily transferable, paving the way for the study of a broad spectrum of other quasiparticles.
Acknowledgements.
W. Li and Y. Lu contributed equally to this work. This work was supported by the National Natural Science Foundation of China under Grants 12374054 and 12022415.